river accelerates a 200 kg motorcycle which results in a force being exerted that  speeds up the motorcycle from 15m/s to 45m/s over a time interval of 13s.


1
Expert's answer
2021-03-14T19:21:23-0400

The force acting on the motorcycle


"F=ma=m\frac{v-v_0}{t}=200\cdot\frac{45-15}{13}\approx462\ (N)" . Answer